Marcus123321:

Find the distance between the points (7,5) and (10,1).

ihatelifeFS420:

do you know the distance formula for that

ihatelifeFS420:

so this is the distance formula d=√(x2-x1)² + (y2-y1)²

ihatelifeFS420:

so do 7-10 and then do 5-1 then add

jhonyy9:

@ihatelifefs420 wrote:
do you know the distance formula for that
perfect help - congrats !

ihatelifeFS420:

@jhonyy9 wrote:
@ihatelifefs420 wrote:
do you know the distance formula for that
perfect help - congrats !
thanks

ihatelifeFS420:

so 7-10=-3 and 5-1=4 and -3+4=1

ihatelifeFS420:

so the distance is 1

Joe348:

It's not 1

Joe348:

Joe348:

This would be the proper way to solve this question.

ihatelifeFS420:

sorry my bad I forgot to do the square root

Joe348:

It happens lol

Spectrum:

Can you close this post? Clears out the posts so we know who has been helped and who needs help.

lilbart:

You can use algebraic calculators like mathway